OpenStudy (anonymous):

2-sin^2x/cos2x=sec^2x+1

OpenStudy (anonymous):

\[ 2 -\frac {\sin^2(x)}{\cos^2(x)}= \sec^2(x) +1\\ 2 -\tan^2(x) = 1+ \tan^2(x) +1\\ 2\tan^2(x)=0\\ \tan(x)=0\\ x= k \pi,\,\, k=0,\pm 1 ,\pm 2 ,\pm 3, \cdots \]